OpenStudy (meredith17):

help? The length of a rectangle is 3 times the width, and the perimeter is 22. Find the dimensions of the rectangle.

OpenStudy (mathstudent55):

Are you studying systems of equations?

OpenStudy (meredith17):

"number problems"

OpenStudy (mathstudent55):

Let L = length Let W = width Using L and W, can you write an equation that represents this sentence? "The length of a rectangle is 3 times the width"

OpenStudy (lapinblanc):

Length is 3 times the width L=3w Perimeter=22 Perimeter formula is L+L+w+w We replace our L with the above 3w and we know perimeter is 22: So 3w+3w+w+w=22 8w=22 Divide out the 8 by both sides w=2.75 We reuse L=3w to find L L=3*w L=3*2.75 L=8.25 So w=2.75 L=8.25 to confirm, you can put it back into perimeter formula to make sure equals 22 2.75+2.75+8.25+8.25=22

OpenStudy (meredith17):

thanks both of you!!!

OpenStudy (meredith17):

i filled it all back in and got 22